### v4

@@ Abstract
In this paper, we consider a class of nonconvex and nonsmooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms. We show the second-order optimality conditions only depend on the nonzeros of the stationary points. We propose two damped iterative reweighted algorithms, including the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems. For DIRL1, we show the reweighted $\ell_{1}$ subproblem has support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point. For DIRL2, we show the solution map of the reweighted $\ell_{2}$ subproblem is differentiable and Lipschitz continuous everywhere. Therefore, the map of DIRL1 and DIRL2 and their inverse are Lipschitz continuous, and the strict saddle points are their unstable fixed points. By applying the stable manifold theorem, these algorithms are shown to converge only to local minimizers with random initialization when the strictly saddle point property is assumed.
@@
strict saddle, iterative reweighted $\ell_{1}$ algorithm, nonconvex regularization, center stable manifold theorem
@@ 1 Introduction
In this paper, we consider the following nonsmooth regularization problem,
$\min\limits_{x\in\mathbb{R}^{n}}F(x):=f(x)+\lambda\psi(x),$ (P)
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$. The data fitting function $f$ is differentiable twice continuously and possibly nonconvex, and $r$ satisfies the following properties.
@@
$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$, C2 smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
@@
The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance. This makes it well-suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [1, 2, 3, 4]. Many widely used nonconvex regularization such as the EXP penalty [38] or the LOG penalty [39], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability. We provide the explicit form of several commonly used nonconvex regularizations in Table 1. It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) nonLipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as the $\ell_{p}$ quasi-norm (LPN).
@@
Numerous studies focus on gradient-based optimization methods that are guaranteed to find first-order stationary points. However, the nonconvexity of $F$ can lead to multiple local minima and saddle points, which can cause algorithms to converge to stationary points rather than the (local) optimum. Additionally, the presence of saddle points can significantly slow down the convergence of optimization algorithms, especially gradient-based methods. At a saddle point, the gradient provides misleading information about the descent directions. As a result, avoiding saddle points can help optimization algorithms more effectively converge to better solutions. Therefore, it becomes a prevalent topic, particularly in high-dimensional optimization problems, and is also the central focus of this paper.
@@
In smooth optimization, many methods have been developed and analyzed to escape saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19]. In smooth settings, it is widely assumed that strict saddle property holds, which means that all saddle points have at least one strictly negative eigenvalue. This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases, and it has been shown to be generally satisfied for smooth problems [20, 21].
@@
In nonsmooth optimization, avoiding saddle points poses a greater challenge. At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate avoidance of saddle points. Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization. An example of this complexity is demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [22]. Recent research [23] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point. Additionally, several proximal methods, including the proximal point method, the proximal gradient method, and the proximal linear method [23], as well as a gradient method incorporating random perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings. In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$ weakly convex, which means that $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. They assumed an active manifold around the limit point and then defined the strict saddle points as those of the $f$ restricted to the manifold. The weakly $\rho$-convexity for $r$ plays a crucial role in constructing the C2 differentiable Moreau envelope with the same stationary points and strict saddle points in the proximal methods.
@@
As for the regularized problem (P), several first-order methods have been proposed. [25] introduces a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1. This method is subsequently extended to address the $\ell_{p,q}$ group norm regularized problems of the group $\ell_{p,q}$ in [26]. Furthermore, the Difference of Convex (DC) algorithm is enlisted [27] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions. It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold true for the $\ell_{p}$ regularized problem. Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularizers. Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [28] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems [29]. Subsequent work in [30] established its global convergence and demonstrated its locally stable support property [31]. Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [32] and its extrapolated version was proposed in [33]. While these first-order methods primarily focus on identifying first-order stationary points, none of them guarantee the avoidance of saddle points.
@@ 1.1 The main results
In this paper, our main purpose is to develop and analyze first-order methods that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity. In particular, we focus on iteratively reweighted methods which approximate the regularization term by simpler weighted $\ell_{1}$ or $\ell_{2}$ norms [34, 35, 36, 30, 37, 29, 31, 33]. In contrast to the proximal gradient algorithm, iteratively reweighted methods solve a sequence of weighted $\ell_{1}$ or $\ell_{2}$ regularized subproblems and are widely applicable to various nonconvex regularization problems.
The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator. This method offers a favorable local property [31], known as the model/support identification, under the following additional assumption.
@@
For any stationary point $x^{*}$, $0\in{\rm rint}\partial F(x^{*})$ (where $\partial F$ denotes the regular subdifferential and ${\rm rint}$ denotes the relative interior).
This kind of assumption is also widely made to guarantee that the existence of an active manifold around a stationary point as assumed by [23]. At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0. In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{I}$ as follows after a finite number of iterations.
Definition 1.
(Model Subspace) The model subspace defined by $\mathcal{I}\subset\{1,\ldots,n\}$ is
:=$\mathcal{M}_{\mathcal{I}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{I}^{c},u_{j}=0\}.$
This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{I}}$. This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) in $\mathcal{M}_{\mathcal{I}}$, although the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
The iterative reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these nonconvex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the solution is trivially given. Compared to ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property. In fact, it can be often witnessed that this algorithm can generate iterates with elements asymptotically approaching zero. To show the different behaviors of the two algorithms, we provide the following example.
Example 1.
Consider the $\ell_{1/2}^{1/2}$ regularization problem
$\min_{x\in\mathbb{R}^{2}}F(x)=f(x)+\|x\|_{1/2}^{1/2},\ \text{ with }f(x)=x_{1}^{2}+(x_{2}-5/4)^{2}.$
The global minimizer is $(0,1)$, the local minimizer111(0,0)00(0,0)( 0 , 0 ) is always a dummy local minimizer of $\ell_{p}$ regularization problems is $(0,0)$ and the saddle point minimizer is $(0,\frac{3-2\sqrt{2}}{4})\approx(0,0.043)$ which is also a local maximizer on the $x_{2}$-axis.
(a) Contour of $F$.
(b) Contour of $F$ around $(0,0)$.
Let $g(x;x^{k})=\nabla f(x^{k})^{T}(x-x^{k})+\tfrac{1}{2}\|x-x^{k}\|^{2}$. The iterations of the iteratively reweighted algorithms are as follows [29]. We use the ${\rm IRL}_{1}$ algorithm with iteration
$x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{2}\frac{1}{(|x_{1}^{k}|+0.1^{k})^{1/2}}|x_{1}|+\tfrac{1}{2}\frac{1}{(|x_{2}^{k}|+0.1^{k})^{1/2}}|x_{2}|$
and the ${\rm IRL}_{2}$ algorithm with iteration
$x^{k+1}=\arg\min_{x_{1},x_{2}}g(x;x^{k})+\tfrac{1}{4}\frac{1}{((x_{1}^{k})^{2}+0.1^{k})^{3/4}}x_{1}^{2}+\tfrac{1}{4}\frac{1}{((x_{2}^{k})^{2}+0.1^{k})^{3/4}}x_{2}^{2}.$
The trajectories of the two algorithms are depicted in Figure 2.
(a) Trajectory of ${\rm IRL}_{1}$
(b) Trajectory of ${\rm IRL}_{2}$
Figure 2: Trajectories generated by iteratively reweighted methods. ${\rm IRL}_{1}$ eventually falls into the subspace (the $x_{2}$-axis), while ${\rm IRL}_{2}$ has nonzero elements for all iterates.
The behavior of ${\rm IRL}_{2}$ brings another challenge in analyzing the avoidance of strict saddle point. Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures the algorithm reverts to a smooth algorithm in the active manifold in a local neighbourhood of the stationary point. However, there is no answer how a nonsmooth first-order algorithm like ${\rm IRL}_{2}$ would behave in the local region of a strict saddle point.
We now summarize our contribution as follows:
@@
    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{I}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{I}}$.
@@
    We develop damped IRL1 algorithms individually for the two types of regularizer, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning that the next iterate is the combination of a traditional IRL1 iterate and its previous iterate. The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ from becoming $\infty$ at each iteration. At the tail end of the iterations, we show that this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{I}$ on the subspace $\mathcal{M}_{\mathcal{I}}$ and decrease the rest components to 0 at a linear speed. We then apply the center stable manifold theorem and prove that the algorithm can avoid the convergence to strict saddle points.
@@
    We develop damped IRL2 algorithms for nonconvex regularized problems and provide global convergence. Under an additional assumption which is satisfied by most common nonconvex regularizers, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere. We also proved the equivalence between the unstable fixed point of the iteration map and the strict saddle points of the regularization problem. Therefore, the avoidance of strict saddle points is derived. This is the first algorithm shown to have this property without using the active-manifold identification property or the stochastic technique in nonsmooth optimization algorithms.
@@ 1.3 Notation
We denote $\|\cdot\|_{p}$ as the quasi-norm $\ell_{p}$ with $p\in(0,1)$, that is, $\|x\|_{p}=(\sum_{i=1}^{n}|x_{i}|^{p})^{\frac{1}{p}}$. This does not define a proper norm because of the lack of subadditivity. For $x\in\mathbb{R}^{n}$, $x_{i}$ denotes the $i$-th element of $x$. The index set of nonzero and zero components of $x$ is defined as $\mathcal{I}(x)=\{i\mid x_{i}\neq 0\}$, and $\mathcal{J}(x)=\{i\mid x_{i}=0\}$, respectively. Moreover, $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ denotes $\mathcal{I}(x^{*})$ and $\mathcal{J}(x^{*})$ for short. ${\rm diag}(x)$ is the diagonal matrix by placing elements of vector $x$ on the main diagonal. ${\rm sign}(x):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the sign function defined by
$[{\rm sign}(x)]_{i}=\left\{\begin{aligned} &1,\ &x_{i}>0,\\ &-1,\ &x_{i}<0,\\ &0,\ &x_{i}=0.\end{aligned}\right.$ (1)
$DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$. Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}^{*}}\subset\mathbb{R}^{\mathcal{I}^{*}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants. If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}f(x)=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$. Denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively. Given a symmetric matrix $H\in\mathbb{R}^{n\times n}$, let $\lambda_{\min}(H)$ be the smallest eigenvalue of $H$. Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the vector of thresholds $w$ is defined $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$. In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.
Regularizer $\psi(x)$ $r^{\prime}(|x_{i}|)$ $r^{\prime\prime}(|x_{i}|)$
EXP [38] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $pe^{-p|x_{i}|}$ $-\frac{1}{p^{2}}e^{-\frac{|x_{i}|}{p}}{p}$
LOG [39] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$ $-\frac{1}{(|x_{i}|+p)^{2}}$
FRA [40] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$ $-\frac{2p}{(|x_{i}|+p)^{2}}$
LPN [7] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $p|x_{i}|^{p-1}$ $p(p-1)|x_{i}|^{p-2}$
TAN [41] $\sum\limits_{i=1}^{n}\arctan\frac{|x_{i}|}{p}$ $\frac{p}{(|x_{i}|^{2}+p^{2})^{2}}$ $-\frac{2p|x_{i}|}{(|x_{i}|^{2}+p^{2})^{2}}$

### v5

@@ Abstract
This paper considers a class of nonconvex and nonsmooth sparse optimization problems, encompassing most existing nonconvex sparsity-inducing terms. We show that their second-order optimality conditions depend only on the nonzeros of the stationary points. We propose two damped iteratively reweighted algorithms, which are the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems. For DIRL1, we show that the reweighted $\ell_{1}$ subproblem has the support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point. For DIRL2, we show that the solution map of the reweighted $\ell_{2}$ subproblem is differentiable and Lipschitz continuous everywhere. Therefore, the solution maps of DIRL1 and DIRL2 and their inverses are Lipschitz continuous, and the strict saddle points are their unstable fixed points. By applying the stable manifold theorem, these algorithms starting from almost every initial point are shown to converge to local minima when the strict saddle point property is assumed.
@@
strict saddle property, iterative reweighted $\ell_{1}$ algorithm, nonconvex regularization, center stable manifold theorem
{AMS}
49J50, 49M20, 49M37, 65K05, 65K10, 90C26, 90C30
@@ 1 Introduction
In this paper, we consider the following nonconvex and nonsmooth regularization problem,
(P) $\min\limits_{x\in\mathbb{R}^{n}}F(x):=f(x)+\lambda\psi(x),$
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a twice continuously differentiable function and $\psi(x)$ is a penalty/regularization term taking the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$ and $r$ satisfying the following properties.
@@
$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$ and C2 smooth on $(0,+\infty)$, $\lim\limits_{t\rightarrow 0^{+}}r^{\prime}(t)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
@@
The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance of the model. This makes it well suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [17, 31, 35, 47]. Many widely used nonconvex regularizations, such as the EXP penalty [5], the LOG penalty [26], and the $\ell_{p}$ quasi-norm penalty ($10<p<$ 1) [18], satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability. We provide the explicit form of several commonly used nonconvex regularizations in Table 1. It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizers with a bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) nonLipschitz continuous regularizers with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as the $\ell_{p}$ quasi-norm (LPN).
@@
Numerous studies focus on gradient-based optimization methods that are guaranteed to find first-order stationary points. However, the nonconvexity of $F$ can lead to multiple local minima and saddle points, which can cause algorithms to converge to a stationary point rather than a local minimum. In addition, the presence of saddle points can significantly slow the convergence of optimization algorithms, especially gradient-based methods. In this situation the gradient provides misleading information about the descent directions at a saddle point. As a result, avoiding saddle points can help optimization algorithms more effectively converge to better solutions. Therefore, it becomes a prevalent topic, particularly in high-dimensional optimization problems, and is also the central focus of this paper.
@@
In smooth optimization, many methods have been developed and analyzed to escape saddle points, including first-order methods [32, 25, 39], momentum-based methods [45], second-order algorithms [33], the algorithmic frameworks for constrained nonlinear problems [30], stochastic algorithms [19, 13], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [21, 22, 27, 12]. In smooth settings, it is widely assumed that strict saddle property holds, which means that all saddle points have at least one strictly negative eigenvalue. This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases, and it has been shown to be generally satisfied for smooth problems [3, 37].
@@
In nonsmooth optimization, avoiding saddle points poses a greater challenge. At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate avoidance of saddle points. Developing nonsmooth optimization algorithms being capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than that in smooth optimization. An example of this complexity was demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [2]. Recent research [15] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming that $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point. Furthermore, several proximal algorithms with damped methods have been proposed to converge to local minima in nonsmooth optimization settings, including the proximal point algorithm, the proximal gradient algorithm, and the damped proximal linear algorithm [15]. Furthermore, an inexact analog of a stochastically perturbed gradient method applied to the Moreau envelope has also been introduced [14]. In these studies, the nonsmooth function $r(x)$ is assumed to be $\rho$ weakly convex, which means that $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. They assumed an active manifold around the limit point and then defined the strict saddle points as those of the $f$ restricted to the manifold. The weak $\rho$-convexity of $r$ plays a crucial role in constructing a $C^{2}$ differentiable Moreau envelope with the same stationary points and strict saddle points in the proximal methods.
@@
As for the regularized problem (P), several first-order methods have been proposed. [46] introduced a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1. This method was subsequently extended to address the $\ell_{p,q}$ group norm regularized problems with $0<q<1\leq p$ in [20]. Furthermore, the difference of convex (DC) algorithm was enlisted [1] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions. It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold for the $\ell_{p}$ regularized problem. Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularization subproblems. Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [7] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems in [44]. Subsequent work in [28] established its global convergence and demonstrates its locally stable support property [43]. Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [42] and its extrapolated version was proposed in [41]. While these first-order methods primarily focus on identifying first-order stationary points, none of them guarantee the avoidance of strict saddle points.
@@ 1.1 The main results
In this paper, our main purpose is to develop and analyze first-order methods that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weak convexity. In particular, we focus on iteratively reweighted methods which approximate the regularization term by simpler weighted $\ell_{1}$ or $\ell_{2}$ norms [8, 10, 23, 28, 38, 44, 43, 41]. In contrast to the proximal gradient algorithm, the iteratively reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which have an explicit solution characterized by the soft-threshold operator. Therefore, it is widely applicable to various nonconvex regularization problems. This method offers a favorable local property [43], known as the model/support identification, under the following additional assumption.
@@
For any stationary point $x^{*}$, $0\in{\rm rint}\partial F(x^{*})$ (where $\partial F$ denotes the regular subdifferential and ${\rm rint}$ denotes the relative interior.)
This kind of assumption is also needed to guarantee the existence of an active manifold around a stationary point as assumed by [15]. At the tail end of the ${\rm IRL}_{1}$ iteration, the sign of the iterates remains unchanged while the nonzero components are kept uniformly bounded away from 0. In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{I}$ as follows after a finite number of iterations.
Definition 1.1.
The model subspace defined by $\mathcal{I}\subset\{1,\ldots,n\}$ is
:=$\mathcal{M}_{\mathcal{I}}:=\{u\in\mathbb{R}^{n}:u_{j}=0,\forall j\in\mathcal{I}^{c}\}.$
This means that every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{I}}$. This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) in $\mathcal{M}_{\mathcal{I}}$, although the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
The iteratively reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these nonconvex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the analytic solution of each subproblem is trivially given. Compared to ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property. In fact, it can be often witnessed that this algorithm can generate iterates with elements asymptotically approaching zero.
The behavior of ${\rm IRL}_{2}$ brings another challenge in analyzing the avoidance of strict saddle points. Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures that the algorithm reverts to a smooth algorithm on the active manifold in a local neighborhood of the stationary point. However, there is no answer to how a nonsmooth first-order algorithm like ${\rm IRL}_{2}$ would behave in the local region of a strict saddle point.
We now summarize our contributions as follows:
@@
    We show that the local minimum and second-order optimality of $F$ are equivalent to those of $F$ restricted in $\mathcal{M}_{\mathcal{I}}$, and we define the strict saddle points of $F$ as those in $\mathcal{M}_{\mathcal{I}}$.
@@
    We develop damped IRL1 algorithms individually for the two types of regularizers, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning that the next iterate is the combination of a traditional IRL1 iterate and its previous iterate. The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ from becoming $\infty$ at each iteration. At the end of the iterations, we show that this algorithm reverts to a damped gradient descent algorithm in the subspace $\mathcal{M}_{\mathcal{I}}$ and reduces the rest components to 0 at a linear speed. We then apply the center stable manifold theorem and prove that the algorithm can avoid convergence to strict saddle points.
@@
    We develop damped IRL2 algorithms for nonconvex and nonsmooth regularized problems and provide global convergence. Under an additional assumption which is satisfied by the $\ell_{p}$ regularization, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere. We also prove the equivalence between a unstable fixed point of the iteration map and a strict saddle point of the nonconvex regularized problem. Therefore, the avoidance of strict saddle points is derived. This is the first algorithm shown to have this property without using the active-manifold identification property or the stochastic technique in nonsmooth optimization setting.
@@ 1.3 Notation
We denote $\|\cdot\|_{p}$ as the quasi-norm $\ell_{p}$ with $p\in(0,1)$, that is, $\|x\|_{p}=(\sum_{i=1}^{n}|x_{i}|^{p})^{\frac{1}{p}}$. This does not define a proper norm because of the lack of subadditivity. The index sets of nonzero and zero components of $x$ are defined as $\mathcal{I}(x)=\{i\mid x_{i}\neq 0\}$, and $\mathcal{J}(x)=\{i\mid x_{i}=0\}$, respectively. Moreover, $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ denotes $\mathcal{I}(x^{*})$ and $\mathcal{J}(x^{*})$ for short. ${\rm diag}(x)$ is the diagonal matrix by placing elements of vector $x$ on the main diagonal. ${\rm sign}(x):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the sign function defined by $[{\rm sign}(x)]_{i}=1$ if $x_{i}>0$, $[{\rm sign}(x)]_{i}=-1$ if $x_{i}<0$ and $[{\rm sign}(x)]_{i}=0$ if $x_{i}=0$.
$DT(x)$ denotes the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$. Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}}\subset\mathbb{R}^{\mathcal{I}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$ with respect to $\mathcal{I}$, respectively. If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}f(x)=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$. Denote $f_{|\mathcal{I}}$ as the restriction of $f$ in $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$; therefore, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}};0_{\mathcal{J}}]).$ Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with $w$ defined as $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$. In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.
Regularizer $\psi(x)$ $r(|x_{i}|)$ $r^{\prime}(|x_{i}|)$ $r^{\prime\prime}(|x_{i}|)$
EXP [5] $\sum\limits_{i=1}^{n}(1-e^{-p|x_{i}|})$ $1-e^{-p|x_{i}|}$ $pe^{-p|x_{i}|}$ $-p^{2}e^{-p|x_{i}|}$
LOG [26] $\sum\limits_{i=1}^{n}\log(1+p|x_{i}|)$ $\log(1+p|x_{i}|)$ $\frac{p}{1+p|x_{i}|}$ $-\frac{p^{2}}{(1+p|x_{i}|)^{2}}$
FRA [16] $\sum\limits_{i=1}^{n}\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{|x_{i}|}{|x_{i}|+p}$ $\frac{p}{(|x_{i}|+p)^{2}}$ $-\frac{2p}{(|x_{i}|+p)^{3}}$
LPN [18] $\sum\limits_{i=1}^{n}|x_{i}|^{p}$ $|x_{i}|^{p}$ $p|x_{i}|^{p-1}$ $p(p-1)|x_{i}|^{p-2}$
TAN [6] $\sum\limits_{i=1}^{n}\arctan(\frac{|x_{i}|}{p})$ $\arctan(\frac{|x_{i}|}{p})$ $\frac{p}{|x_{i}|^{2}+p^{2}}$ $-\frac{2p|x_{i}|}{(|x_{i}|^{2}+p^{2})^{2}}$